OpenStudy (anonymous):

For a beverage, Lisa mixes cups of lemonade to cups of iced tea in a ratio of 3:7. Which of the following best describes the amount of lemonade to iced tea? For every 3 cups of lemonade, there are 7 cups of iced tea. For every 3 cups of lemonade, there are 4 cups of iced tea. For every 7 cups of lemonade, there are 3 cups of iced tea For every 7 cups of lemonade, there are 4 cups of iced tea. @Dani_Rose

OpenStudy (anonymous):

@Da

OpenStudy (anonymous):

@Dani_Rose

OpenStudy (dani_rose):

I think it's A but i'm not sure

OpenStudy (anonymous):

On a farm, the farmer decides to give pizza to her 15 ducks as a special treat. She orders 3 pizzas, and the total price is $23.46. What is the unit price of each pizza? $1.56 $5.00 $7.82 $70.38

OpenStudy (dani_rose):

23.46/3=7.82

OpenStudy (anonymous):

A biker traveled 24 miles in 3 hours. A unit rate to describe this would be 8 miles per hour. What is another unit rate in this situation? 1 over 8 mile per hour 1 over 8 hour per mile 1 mile per hour 3 hours per mile

OpenStudy (dani_rose):

24/3=8 miles per hour 8=1 hour

OpenStudy (anonymous):

Did you know that chickens sometimes eat vegetables and fruits? A farmer likes to give the chickens apples to share. The farmer observes that 5 chickens eat 8 apples. What is the ratio of the number of chickens to apples? 3:8 5:8 8:3 8:5

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (dani_rose):

3:8

OpenStudy (anonymous):

Two stores offer differently priced sales for the same paper towels, according to the table below. Which store offers the better deal, based on a lower unit price per roll? Store 1 Store 2 Paper towel rolls 15 18 Price $3.60 $3.80 Store ___ has the better deal.

OpenStudy (dani_rose):

Store 2 has the better deal. 3.80/18=0.21 3.60/15=0.24

OpenStudy (anonymous):

which one is the answer?

OpenStudy (dani_rose):

Store 2

OpenStudy (anonymous):

Part A: Use ratio language to describe the relationship of any two things you can think of. Example: For every 5 minutes of jogging, I ran 500 yards. Be creative and come up with your own idea. Part B: Determine the unit rate in your example.

OpenStudy (dani_rose):

I'm not too good with ratios sorry

OpenStudy (misssmartiez):

"mixes cups of lemonade to cups of iced tea " correct on the first one.

OpenStudy (anonymous):

..ok thanks for helping tho.

OpenStudy (anonymous):

Part A: Use ratio language to describe the relationship of any two things you can think of. Example: For every 5 minutes of jogging, I ran 500 yards. Be creative and come up with your own idea. Part B: Determine the unit rate in your example.

OpenStudy (misssmartiez):

I got one. You have 14 blue cards, and 18 red cards.

OpenStudy (misssmartiez):

However, they ask for the unit rate, so we need something that has to do with?

OpenStudy (anonymous):

@hann

OpenStudy (anonymous):

@HannahA

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (misssmartiez):

Lila is a runner. Each

OpenStudy (misssmartiez):

46 MINUTES, she completes 1,000 miles.

OpenStudy (misssmartiez):

or, Amy stopped at Seven 11 for some gas. She paid 30 dollars for 100 gallons.

OpenStudy (hannaha):

Well since unit rate means that we are looking for how much of something inside something else (in simple terms) so if you have "Jill eats 4 apples an hour, for every four apples she eats, Jared eats 6 apples." than you're comparing the amount of apples that jill eats per hour with hw many Jared eats per hour

OpenStudy (anonymous):

Thanks so much guys!.

OpenStudy (misssmartiez):

figure out the unit rate.